Please write in Microsoft Word

  1. Sinclair Oil & Gas, a large energy conglomerate, jointly processes purchased hydrocarbons to generate two intermediate products: ING4, and XGE3. These intermediate products are further processed separately to produce Crude Oil and Natural Gas Liquids (NGL). In June 2019, the following is the monthly data for all the products:

Joint costs

Crude Oil

Natural Gas Liquids

Joint Costs (costs of Hydrocarbons and other inputs and processing to split-off point)

$1,800

Separable costs of processing 400 tons of ING4 into 150 tons of Crude Oil

$175

Separable cost of processing 500 tons of XGE3 into 50 tons of Natural Gas Liquids

$105

Transfer for further processing (tons)

ING4

XGE3

Crude Oil

Natural Gas Liquids

Beginning inventory (tons)

0

0

0

0

Production (tons)

400

500

150

50

Sales (tons)

400

500

150

50

Selling price per ton

$8

$5

$18

$15

  1. Allocate the joint costs of $1,800 between ING4 and XGE3 under
  1. the sales value at split-off method (1 pt)
  2. the physical-measure method. (1pt)
  1. Allocate the joint costs of $1,800 between crude oil and natural gas liquids under the NRV method. (2pts)
  2. Should XGE3 be processed further or sold at the split-off point? (1pt)
Answer #1

a.

Total sales value = 3,200 + 2,500 = $5,700

Sales Value of ING4 = 400 x $8 = $3,200 (3200 / 5700 x 100 = 56.14%)

Sales Value of XGE3 = 500 x $5 = $2,500 (2500 / 5700 x 100 = 43.86%)

Joint cost allocated to ING4 = 1800 x 56.14% = $1010

Joint cost allocated to XGE3 = 1800 x 43.86% = $790

b.

Total Production volume = 900

Production volume (ING4) = 400 (400/900x100 = 44.44%)

Production volume (XGE3) = 500 (500/900x100 = 55.56%)

Joint cost allocated to ING4 = 1800 x 44.44% = $800

Joint cost allocated to XGE3 = 1800 x 43.86% = $1000

c.

NRV = Sales value - Seperable costs

Net Realizable Value - NRV of Crude Oil = ((150 x $18) - $175) = $2525 (79.65%)

Net Realizable Value - NRV of Natural Gas Liquids = ((50 x $15) - $105) = $645 (20.35%)

Total NRV = 2525 + 645 = $3170

Joint cost allocated to ING4 = 1800 x 79.65% = $1434

Joint cost allocated to XGE3 = 1800 x 20.35% = $366

d.

XGE3 Contribution without Processing further = Sales value - Allocated joint cost

= 2500 - 790 = $1710

XGE3 Contribution after Processing further = NRV - Allocated joint cost

= 645 - 366 = $279

XGE3 should not be processes further to Natural Gas Liquids.